A bus moving at a speed of 25m/s was brought to rest in 20 seconds, the retardation is what?
1
Expert's answer
2019-11-11T15:53:16-0500

The retardation is 

"a=\frac{V-v}{t}"

"a=\frac{25-0}{20}=1.25\frac{m}{s^2}"